Amegy Bank can borrow either $20 million or €20 million. The current spot rate of the euro is $1.14. Furthermore, Amegy Bank expects the spot rate of the euro to be $1.10 in 90 days. What is Amegy Bank's dollar profit from speculating if the spot rate of the euro is indeed $1.10 in 90 days? Assume 360 days in a year for simplicity and set your decimal to at least 8 places.

Lending Rate Borrow Rate
U.S. 6.73% 7.20%
Euro 6.80% 7.28%

A. $579,845.

B. $583,800.

C. $588,200.

D. $783,210

E. $980,245.

Please help! Will vote
Answer #1

Following steps will be undertaken:

Borrow Euro 20,000,000

Convert into USD today = 20,000,000*1.14 = $22,800,000

Invest for 90 days and get 22,800,000*(1+6.73%*90/360) = $23,183,610

Convert back to Euros and get 23,183,610/1.10 = Euro 21,076,009.09

Repay Loan 20,000,000*(1+7.28%*90/360) = Euro 20,364,000

Arbitrage Profit = Euro 712,009.09

i.e. $712,009.09*1.10 = $783,209.999

i.e. $783,210

Hence, the answer is D. $783,210